# Stochastic Subspace via Probabilistic Principal Component Analysis for Characterizing Model Error

### v1

@@ 6.2 Parametric linear static problem
To characterize the error induced by ROM, we construct SROMs by replacing the deterministic ROB $\mathbf{V}$ with SROBs $\mathbf{W}$. For the SS-PPCA method, the SROBs $\mathbf{W}$ are sampled using the stochastic subspace model described in Section 3.3 and Algorithm 1. To achieve consistent and sharp UQ, the hyperparameter $\beta$ is optimized by minimizing the objective function given in Eq. 8. For this example, $\mathbf{u}_{\_}E$ represents the HDM displacement at training parameters, and $\mathbf{u}_{\_}L^{o}$ represents the ROM displacement at the same parameters. The optimal value of the hyperparameter $\beta$ for the SS-PPCA method is 20, which is optimized efficiently using the strategy given in Section 4. The results for the NPM method are generated following the strategy from [32]. Since the RSM method requires multiple models for its application, we construct five models as ROMs with 4-dim bases $\mathbf{V}^{(l)}=\mathbf{\Phi}^{(l)}\mathbf{Q}^{(l)}$, where $\mathbf{\Phi}^{(l)}=[\boldsymbol{\phi}_{\_}i]_{\_}{i=2,\cdots,6}^{i\neq l}$, $l\in\{2,\cdots,6\}$, and $\mathbf{Q}^{(l)}=\pi(\mathbf{\Phi}^{(l)\intercal}\mathbf{V})\in O(k)$ solves the Procrustes problem $\min_{\_}{\mathbf{Q}\in O(k)}\|\mathbf{V}-\mathbf{\Phi}^{(l)}\mathbf{Q}\|_{\_}{\text{F}}$. These models are designed so that the model subspaces are distinct invariant subspaces of $\mathbf{K}$, with bases best aligned with the global basis $\mathbf{V}$ in the Frobenius norm. We then apply the strategy described in [37] to characterize ROM error.
Figure 4: HDM vs ROM displacement at the test parameter
Figure 5: Difference in ROM and SROM mean displacement
@@
Figure 7: Linear static problem: prediction by NPM.
Figure 8: Linear static problem: prediction by RSM.
From Figure 5, we can observe that all three methods are broadly comparable regarding the difference in the ROM and SROM mean displacements. However, it is important to note that the NPM method has steep, short-range jumps in the error near the boundaries. This behavior is due to the explicit implementation of boundary conditions in the NPM method. In contrast, SS-PPCA and RSM methods satisfy boundary conditions by default.
Figures 6, 7 and 8 shows the ROM error characterization for the linear parametric system at the test parameters. All three methods yield consistent UQ, with the SS-PPCA result being slightly sharper. Note that some hyperparameters of the NPM method, such as the weight and the scale, are not optimized. Therefore, we must use the trial-and-error approach to achieve the optimal result shown in Figure 7. This makes the NPM method computationally intensive and not fully automated. It roughly takes 3.33 hours for the NPM method, whereas the SS-PPCA method only takes 9.1 seconds, significantly saving computational resources.
@@ 6.3 Static problem: characterizing HDM error
To characterize both the ROM and HDM error, we construct SROM by substituting the deterministic ROB $\mathbf{V}$ with the SROB $\mathbf{W}$. For the SS-PPCA method, the SROBs $\mathbf{W}$ are sampled using the stochastic subspace model described in Section 3.3 and Algorithm 1. To have consistent and sharp UQ, the hyperparameter $\beta$ is optimized by minimizing the objective function given in Eq. 8. For this example, $\mathbf{u}_{\_}E$ represents the sparsely observable experimental displacement, and $\mathbf{u}_{\_}L^{o}$ denotes the ROM displacement at the same locations. The optimum value of the hyperparameter $\beta$ for the SS-PPCA method comes out to be 8, which is optimized efficiently using the strategy given in Section 4. The results for the NPM method are generated following the strategy from [32]. Since there is only one model in this example, we exclude the RSM method from the comparison.
@@
Figure 10: Linear static problem with experimental data: NPM prediction.
Figures 9 and 10 shows the HDM error characterization by the SS-PPCA and NPM methods using the sparse experimental observations. Both approaches successfully capture the model discrepancy, yielding accurate uncertainty quantification for the HDM predictions. Meanwhile, the SS-PPCA method provides a slightly sharper UQ than the NPM method. The more significant difference between the two methods lies in their computational cost, particularly during hyperparameter optimization. SS-PPCA involves only a single hyperparameter $\beta$, which is optimized efficiently using the training procedure outlined in Section 4. In comparison, the NPM method involves a large number of hyperparameters, and the optimization process is computationally challenging. In this experiment, NPM training is approximately 5,440 times slower than that of SS-PPCA, see Table 2. Moreover, the NPM method does not have a systematic approach for selecting its scale hyperparameter. As a result, we run the method four times with different scaling factors to determine the most suitable value empirically. This trial-and-error strategy requires user intervention and reduces efficiency. In contrast, SS-PPCA offers a fully automated, efficient, and principled training workflow with minimal user input. This makes it well suited for practical applications.
@@ 6.4 Dynamics problem: space structure
In Sections 6.2 and 6.3, we demonstrated that both the NPM and SS-PPCA methods perform well in characterizing uncertainty despite the notably higher training time of the NPM method. However, those cases involved relatively simple linear static problems. To evaluate the robustness and scalability of these methods, we compare them in a much more complex linear dynamics problem of a space structure. Because of the higher model complexity and more realistic structural dynamics, this example may better assess the efficiency and effectiveness of the two methods.
@@
where reduced matrices are denoted with a pattern $\mathbf{A}_{\_}{\mathbf{V}}:=\mathbf{V}^{\intercal}\mathbf{A}\mathbf{V}$, so that $\mathbf{M}_{\_}{\text{H},\mathbf{V}}=\mathbf{V}^{\intercal}\mathbf{M}_{\_}{\text{H}}\mathbf{V}$, for example. The initial condition of ROM is given by $\mathbf{q}(0)=0$ and $\dot{\mathbf{q}}(0)=0$. In contrast to the high simulation time of HDM, ROM takes 0.2 seconds—roughly 11,200 times faster. However, this computational efficiency comes at the expense of reduced accuracy. Hence, error characterization of ROM is essential, which is done by SROM throughout this study. To construct the SROM, we substitute the deterministic ROB $\mathbf{V}$ by the SROB $\mathbf{W}$.
For the SS-PPCA method, the SROBs $\mathbf{W}$ are sampled using the stochastic subspace model described in Section 3.3 and Algorithm 1. For this example, $\mathbf{u}_{\_}E$ corresponds to the velocity from the HDM at a critical structural node, and $\mathbf{u}_{\_}{L}^{o}$ corresponds to the velocity predicted by ROM at the same location. The optimum value of the hyperparameter $\beta$ for the SS-PPCA method comes out to be 39, which is optimized using the strategy given in Section 4. For comparison, we include the NPM method results obtained by following the procedure of [32].
The HDM velocity shows a high-frequency behavior, especially during the initial 10 milliseconds. The ROM with just 10 modes fails to capture this behavior. Figures 12 and 13 compares the velocity prediction of SS-PPCA and NPM. It can be observed that the 95% PI of the SS-PPCA method is notably consistent and significantly sharper than that of the NPM. The NPM method does not capture the high-frequency behavior of HDM velocity, leading to underestimation (see the Figure 13 inset for $[0,10]$ ms). In contrast, the SS-PPCA method not only captures it but also maintains tighter bounds around it. Additionally, the mean SROM velocity of SS-PPCA closely matches the deterministic ROM velocity, as it should be. This behavior is not observed in the NPM results, where the mean prediction deviates from the ROM.
Figures 14 and 15 show the $x$-direction acceleration at the critical node predicted by the SS-PPCA and the NPM methods. The acceleration also exhibits a high-frequency behavior, due to the impulse loading. Despite this, our SS-PPCA method is able to produce consistent UQ with sharper bounds than the NPM method. It is important to note that even though velocity (the QoI for this example) is used for training the hyperparameters and displacement is used for obtaining the deterministic ROB $\mathbf{V}$, our approach enables predictions of unobserved quantities such as acceleration. The fact that acceleration has not been observed makes our result very intriguing. Furthermore, our approach enables efficient predictions of any unobserved QoIs of the whole system, a capability that many studies in the literature fail to achieve. This highlights the effectiveness and practicality of our approach to engineering applications.
Figures 16 and 17 compares the $x$-direction displacement at the critical node predicted by the SS-PPCA and the NPM methods. It can be observed that SS-PPCA displacement prediction is highly accurate: the SROM, ROM, and HDM show similar behavior, and the 95% PI is very tight. In contrast, the NPM method does not provide a sharp prediction and underestimates the displacement (see the Figure 17 inset for $[68,76]$ ms). One reason for this may be that the global optimization of NPM hyperparameters is highly challenging. In contrast, the SS-PPCA method, with only one hyperparameter, is much easier to train.
@@
Figure 13: Dynamic prediction of the NPM model: velocity.
@@
Figure 17: Dynamic prediction of the NPM model: displacement.
@@
All the quantities compared above, whether observed or unobserved, are from the same node of the space structure. However, we may need to make a prediction at a different node of the system, or we may be interested in the general behavior of the system. Figure 18 shows the prediction of the SS-PPCA method for the $y$-direction velocity at a random node; the result is still consistent and very sharp. This shows that the SS-PPCA methods allow us to make predictions about the entire structure just by training on the observation data of a single DoF.
Table 1 quantifies the coverage percentage and average width of the 95% PIs after the initial 10 ms. While both methods are mostly consistent, SS-PPCA is at least a few times sharper than NPM, depending on the QoI being displacement, velocity, or acceleration.
Table 1: Dynamics problem: consistency and sharpness of 95% PI.
SS-PPCA NPM
95% PI coverage width coverage width width ratio*
$d_{\_}x$ ($10^{-4}$ in) 96.78% 3.26 100.0% 102.8 45
$v_{\_}x$ (in/s) 94.83% 1.22 96.7% 3.62 2.92
$a_{\_}x$ ($10^{3}$ in/s2) 85.49% 7.433 89.15% 11.94 1.74
    Average ratio of 95% PI widths: NPM vs. SS-PPCA.
6.5 Computational costs and discussion
Table 2: Computational cost analysis: hyperparameter optimization
Computational cost (in secs) SS-PPCA NPM RSM Ratio*
Ex1 2.5 34.2 0.12 13.7
Ex2 9.1 $12,020$ 0.34 1320.9
Ex3 1.9 $10,350$ NA $5,447.4$
Ex4 785 $208,970$ NA $266.2$
    Computational cost ratio: NPM vs. SS-PPCA.
Table 2 compares the computational cost of hyperparameter optimization for NPM, RSM, and SS-PPCA across all four numerical examples. Although the RSM method uses the least time of all three methods, it can only be applied in a multiple-model case. Therefore, we mainly focus on comparing the computational cost of NPM and SS-PPCA. It is evident that SS-PPCA requires significantly less time for hyperparameter optimization than NPM. This difference is primarily due to the number of hyperparameters in each method. The number of hyperparameters in the NPM method is proportional to the square of the ROM dimension, whereas SS-PPCA has only one. The NPM method was further modified in a recent study [39], where the authors attempted to reduce the number of hyperparameters by using an autoencoder. Additionally, the authors sought to reduce the computational time by computing analytical gradients of the objective function with respect to the hyperparameters. However, the method is challenging to implement. Hence, we compare SS-PPCA performance with that of the original NPM method [32]. One issue with methods based on SROM is that they cannot fully eliminate model error as they rely on a ROM. This issue can be addressed by increasing the number of reduced bases or incorporating model closure terms. A study on model closure for ROMs is presented by Ahmed et al. [45].
@@ 7 Conclusion
We introduced a novel stochastic subspace model, which is used to characterize model error. It is simple, efficient, easy to implement, and well-supported by analytical results. Through various numerical examples, we reveal the characteristics of this model, show its flexibility across single and multiple model cases, establish its consistency, and quantify its remarkable improvement in sharpness over current methods. It opens up a promising path to the challenging problem of model-form uncertainty.
Unlike current approaches, our SS-PPCA model does not have any hyperparameter besides a concentration parameter that controls scale, which is optimized systematically. This design not only improves consistency but also eliminates hand tuning.

### v2

@@ 6.1 Parametric linear static problem
To characterize the error induced by ROM, we construct SROMs by replacing the deterministic ROB $\mathbf{V}$ with SROBs $\mathbf{W}$. The SROBs $\mathbf{W}$ are sampled using the stochastic subspace model described in Section 3.3 and Algorithm 1. To achieve consistent UQ, the hyperparameter $\beta$ is optimized by minimizing the objective function given in Eq. 8. For this example, $\mathbf{u}_{\_}E$ represents the HDM displacement at training parameters, and $\mathbf{u}_{\_}L^{o}$ represents the ROM displacement at the same parameters. The optimal value of the hyperparameter $\beta$ is 20, which is optimized efficiently using the strategy given in Section 4. The hyperparameter training time for this example is 9.1 seconds.
@@
Figure 3 shows the ROM error characterization for the linear parametric system at the test parameters. It can be observed that the SROM mean displacement is very close to the ROM displacement, and the SROM 95% PI is able to capture the ROM error consistently and sharply. The coverage for this example is 99%, that is, the SROM 95% PI has captured the ROM error over most of the domain. It can also be observed that this coverage is achieved while maintaining tight bounds around the HDM displacement.
@@ 6.2 Static problem: characterizing HDM error
To characterize both the ROM and HDM error, we construct SROMs by substituting the deterministic ROB $\mathbf{V}$ with the SROBs $\mathbf{W}$. The SROBs $\mathbf{W}$ are sampled using the stochastic subspace model described in Section 3.3 and Algorithm 1. To have consistent UQ, the hyperparameter $\beta$ is optimized by minimizing the objective function given in Eq. 8. For this example, $\mathbf{u}_{\_}E$ represents the sparsely observable experimental displacement, and $\mathbf{u}_{\_}L^{o}$ denotes the ROM displacement at the same locations. The optimum value of the hyperparameter $\beta$ is 8, which is optimized efficiently using the strategy given in Section 4. The hyperparameter training time for this example is 1.9 seconds.
@@
Figure 4 shows the HDM error characterization by the SS-PPCA method using the sparse experimental observations. It can be observed that our method successfully captures the model discrepancy, yielding accurate uncertainty quantification for the HDM predictions. The SROM 95% PI captures the noiseless ground-truth data well, and the bounds are very sharp. The trained SROM model via SS-PPCA enables fast prediction with error estimates and eliminates the dependence on computationally expensive HDM.
@@ 6.3 Dynamics problem: space structure
In Sections 6.1 and 6.2, we demonstrated that the SS-PPCA method performs well in characterizing uncertainty with low computational cost. However, those cases involved relatively simple linear static problems. To evaluate the robustness and scalability of the method, we compare it in a much more complex linear dynamics problem of a space structure. Because of the higher model complexity and more realistic structural dynamics, this example may better assess the efficiency and effectiveness of the method.
@@
where reduced matrices are denoted with a pattern $\mathbf{A}_{\_}{\mathbf{V}}:=\mathbf{V}^{\intercal}\mathbf{A}\mathbf{V}$, so that $\mathbf{M}_{\_}{\text{H},\mathbf{V}}=\mathbf{V}^{\intercal}\mathbf{M}_{\_}{\text{H}}\mathbf{V}$, for example. The initial condition of ROM is given by $\mathbf{q}(0)=0$ and $\dot{\mathbf{q}}(0)=0$. In contrast to the high simulation time of HDM, ROM takes 0.2 seconds—roughly 11,200 times faster. However, this computational efficiency comes at the expense of reduced accuracy. Hence, error characterization of ROM is essential, which is done by SROM throughout this study. To construct the SROMs, we substitute the deterministic ROB $\mathbf{V}$ by the SROBs $\mathbf{W}$.
For the SS-PPCA method, the SROBs $\mathbf{W}$ are sampled using the stochastic subspace model described in Section 3.3 and Algorithm 1. For this example, $\mathbf{u}_{\_}E$ corresponds to the velocity from the HDM at a critical structural node, and $\mathbf{u}_{\_}{L}^{o}$ corresponds to the velocity predicted by ROM at the same location. The optimum value of the hyperparameter $\beta$ comes out to be 39, which is optimized using the strategy given in Section 4. The hyperparameter training time for this example is 785 seconds, which is low compared to the computational cost of the HDM. This shows that our method is scalable and can provide fast and reliable predictions even in complex systems.
@@
The HDM velocity shows a high-frequency behavior, especially during the initial 10 milliseconds. The ROM with just 10 modes fails to capture this behavior. Figure 6 illustrates that the 95% PI of the SS-PPCA method is notably consistent and significantly sharp, even in the initial 10 milliseconds period (see the Figure 6 inset for $[0,10]$ ms). The SS-PPCA method not only captures the high-frequency behavior but also maintains tight bounds around it. Additionally, the SROM mean velocity of SS-PPCA closely matches the deterministic ROM velocity, as it should be. The coverage for velocity via the SS-PPCA method is 94.83%.
@@
Figure 7 shows the $x$-direction acceleration at the critical node predicted by the SS-PPCA. The acceleration also exhibits a high-frequency behavior, due to the impulse loading. Despite this, our SS-PPCA method is able to produce consistent UQ with sharp bounds. It is important to note that even though velocity (the QoI for this example) is used for training the hyperparameters and displacement is used for obtaining the deterministic ROB $\mathbf{V}$, our approach enables predictions of unobserved quantities such as acceleration. The fact that acceleration has not been observed makes our result very intriguing. Furthermore, our approach enables efficient predictions of any unobserved QoIs of the whole system, a capability that many studies in the literature fail to achieve. This highlights the effectiveness and practicality of our approach to engineering applications. The coverage for acceleration via the SS-PPCA method is 85.49%, which means that the error characterization can be improved a bit. However, given that the acceleration is an unobserved quantity and exhibits a high-frequency behavior, the result is quite good.
Figure 8 compares the $x$-direction displacement at the critical node predicted by the SS-PPCA. It can be observed that SS-PPCA displacement prediction is highly accurate: the SROM, ROM, and HDM show similar behavior, and the 95% PI is very tight. The coverage for displacement via the SS-PPCA method is 96.78 %, which means that the displacement prediction is consistent and highly accurate.
@@
All the quantities compared in Figures 6, 7 and 8, whether observed or unobserved, are from the same node of the space structure. However, we may need to make a prediction at a different node of the system, or we may be interested in the general behavior of the system. Figure 9 shows the prediction of the SS-PPCA method for the $y$-direction velocity at a random node; the result is still consistent and very sharp. This shows that the SS-PPCA methods allow us to make predictions about the entire structure just by training on the observation data of a single DoF.
In summary, the low computational cost of hyperparameter optimization of SS-PPCA, combined with the low computational cost of ROM, can accelerate the prediction of complex engineering phenomena. Furthermore, SS-PPCA offers a fully automated, efficient, and principled training workflow with minimal user input. This makes it well suited for practical applications. One issue with methods based on SROM is that they cannot fully eliminate model error as they rely on a ROM. This issue can be addressed by increasing the number of reduced bases or incorporating model closure terms. A study on model closure for ROMs is presented by Ahmed et al. [46].
@@ 7 Conclusion
We introduced a novel stochastic subspace model, which is used to characterize model error in the framework of stochastic reduced-order models. It is simple, efficient, easy to implement, and well-supported by analytical results. Our SS-PPCA model has only one hyperparameter, which is optimized systematically to improve consistency. Through various numerical examples, we reveal the characteristics of this model, show its flexibility across single and multiple model cases, establish its consistency, and quantify its remarkable sharpness. It opens up a promising path to the challenging problem of model-form uncertainty.